how do you solve :
1-2(e-1)=-3-4(3-5)
the answer should be e=7
but i'm just not getting it.
Can you explain step by step please

Found 2 solutions by scott8148, solver91311:
Answer by scott8148(6628)

distributing the -2 and the -4 ___ 1 - 2e + 2 = -3 - 12 + 20

consolidating terms ___ -2e + 3 = 5

subtracting 3 ___ -2e = 2

dividing by -2 ___ e = -1

sorry , no 7...

Answer by solver91311(24713)



Nope.



If you replace with 7, you get:



Clear the parentheses:



Do the multiplications:



Clear the parentheses:



Do the addition:



So, start again:



Calculate the value of the RHS. It is the same value we obtained doing the work already illustrated, namely 5:



Add -1 to both sides:



Divide both sides by -2:



Add 1 to both sides



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism